### v3

# GSL-LPA: Fast Label Propagation Algorithm (LPA) for Community Detection with no Internally-Disconnected Communities

Abstract.
Community detection is the problem of identifying tightly connected clusters of nodes within a network. Efficient parallel algorithms for this play a crucial role in various applications, especially as datasets expand to significant sizes. The Label Propagation Algorithm (LPA) is commonly employed for this purpose due to its ease of parallelization, rapid execution, and scalability — however, it may yield internally disconnected communities. This technical report introduces GSL-LPA, derived from our parallelization of LPA, namely GVE-LPA. Our experiments on a system with two 16-core Intel Xeon Gold 6226R processors show that GSL-LPA not only mitigates this issue but also surpasses FLPA, igraph LPA, and NetworKit LPA by $\times5$, 10,300×10,300$\times$10 , 300 ×, and 5.8×5.8$\times$5.8 ×, respectively, achieving a processing rate of $844M$ edges/s on a $3.8B$ edge graph. Additionally, GSL-LPA scales at a rate of 1.6×1.6$\times$1.6 × for every doubling of threads.
Community detection, Internally-disconnected communities, Parallel Label Propagation Algorithm (LPA)
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NY
1. Introduction
Community detection refers to the task of identifying groups of vertices characterized by dense internal connections and sparse connections between groups (Fortunato, 2010). These groups, often referred to as communities or clusters (Abbe, 2018), offer valuable insights into the structure and function of networks (Fortunato, 2010). Community detection has a wide range of applications across multiple fields. In ecology, it is applied to assess if food webs are structured into compartments, where species within each compartment interact more frequently with each other than with those in different compartments (Guimerà et al., 2010). Within healthcare, it helps in analyzing the dynamics of populations vulnerable to epidemic diseases (Salathé and Jones, 2010), detecting conditions such as lung cancer (Bechtel et al., 2005), and classifying tumor types using genomic data (Haq and Wang, 2016). In drug discovery, community detection aids in identifying groups of similar compounds or target proteins, which supports the discovery of novel therapeutic agents (Ma et al., 2019). Furthermore, community detection has been effectively used to reveal the structure and evolution of complex biological networks, including metabolic networks (Kim et al., 2009), Gene Regulatory Networks (GRNs) (Rivera et al., 2010), and Lateral Gene Transfer (LGT) networks (Popa et al., 2011). It also plays a significant role in understanding human brain networks, enabling insights into their structural and functional organization (He and Evans, 2010).
The problem of community detection is due to the lack of prior knowledge about the number of communities and their size distribution (Gregory, 2010). Heuristic-based approaches are commonly employed for community detection (Raghavan et al., 2007; Blondel et al., 2008; Xie et al., 2011; Traag and Šubelj, 2023; Clauset et al., 2004; Duch and Arenas, 2005; Reichardt and Bornholdt, 2006; Kloster and Gleich, 2014; Traag et al., 2019; You et al., 2020; Rosvall and Bergstrom, 2008; Whang et al., 2013). The modularity metric (Newman, 2006a) is used to measure the quality of communities identified. The communities identified are considered intrinsic when based solely on network topology and disjoint when each vertex belongs to only one community (Gregory, 2010; Coscia et al., 2011). The Label Propagation Algorithm (LPA), also known as RAK (Raghavan et al., 2007), is a widely adopted diffusion-based heuristic for community detection, capable of identifying communities of moderate quality. It offers simplicity, speed, and scalability advantages over the Louvain method (Blondel et al., 2008), another effective community detection algorithm renowned for its high-quality results. Specifically, we observe LPA to outpace Louvain by 2.314×2.3-14$\times$2.3 - 14 × in terms of speed, while identifying communities of $3-30\%$ lower quality. LPA’s efficiency is attributed to its avoidance of repeated optimization steps and its ease of parallelization. Thus, LPA is well-suited for applications prioritizing high performance at the expense of slightly lower result quality. While LPA typically yields communities with lower modularity scores, it has been observed to achieve high Normalized Mutual Information (NMI) score relative to the ground truth (Peng et al., 2014). In our experimentation with alternative label-propagation based methods such as COPRA (Gregory, 2010), SLPA (Xie et al., 2011), and LabelRank (Xie et al., 2013), we found LPA to be the most efficient, while identifying communities of equivalent quality (Sahu, 2023d).
However, LPA is susceptible to identifying a large number of internally disconnected communities. Through our experimental analysis, we observe that up to $3.1\%$, $0.3\%$, and $14.5\%$ of communities identified using FLPA (Traag and Šubelj, 2023), igraph LPA (Csardi et al., 2006), and NetworKit LPA (Staudt et al., 2016) exhibit this issue. Rectifying such disconnected communities is crucial to ensuring the accuracy and robustness of community detection algorithms. While a number of studies have proposed addressing internally-disconnected communities as a post-processing step, they do not present a concrete algorithm to do so (Raghavan et al., 2007; Gregory, 2010; Hafez et al., 2014; Luecken, 2016; Wolf et al., 2019). In this technical report, we present a GSL-LPA, which is based upon an efficient multicore implementation of LPA (Sahu, 2023c), and utilizes our BFS-based approach for splitting internally-disconnected communities — which we show, outperforms other approaches for splitting the identified disconnected communities.
2. Related work
A number of schemes have been devised for community detection (Raghavan et al., 2007; Blondel et al., 2008; Xie et al., 2011; Traag and Šubelj, 2023; Clauset et al., 2004; Duch and Arenas, 2005; Reichardt and Bornholdt, 2006; Rosvall and Bergstrom, 2008; Whang et al., 2013; Kloster and Gleich, 2014; Traag et al., 2019; You et al., 2020). These can be categorized into three basic approaches: bottom-up, top-down, and data-structure based (Souravlas et al., 2021). Additionally, they can be classified as divisive, agglomerative, or multi-level methods (Zarayeneh and Kalyanaraman, 2021). To evaluate these methods, fitness scores like modularity (Newman, 2006b) are used.
The Label Propagation Algorithm (LPA) is a diffusion-based method for identifying communities. Techniques for parallelizing LPA include vertex assignment with guided scheduling (Staudt and Meyerhenke, 2015), parallel bitonic sort (Soman and Narang, 2011), and pre-partitioning of the graph (Kuzmin et al., 2015). In terms of quality, LPA may yield communities with low modularity score, resulting in the emergence of a dominant community that envelops the majority of the nodes, obscuring finer community structures (Gregory, 2010). To address instability of result, community structure quality, and performance due to vertex processing order, a few ordering strategies have been attempted. These include updating only the active nodes (Xie and Szymanski, 2011) or prioritizing specific subsets based on properties, like core or boundary nodes (Gui et al., 2018). In some cases, oscillations in label assignments may arise instead of convergence. To address this, asynchronous mode (Leung et al., 2009), alternating updates of independent node subsets (Cordasco and Gargano, 2012), and parallel graph coloring techniques (Cordasco and Gargano, 2012) have been attempted. Raghavan et al. (Raghavan et al., 2007) note that after just five iterations, labels of $95\%$ of nodes converge to their final values.
Additional enhancements to LPA encompass employing a stable (non-random) mechanism for selecting labels when encountering multiple best labels (Xing et al., 2014), mitigating the challenge posed by monster communities (Berahmand and Bouyer, 2018; Sattari and Zamanifar, 2018), discerning central nodes and combining communities to enhance modularity (You et al., 2020), as well as leveraging frontiers with alternating push-pull to minimize the number of edges traversed and enhance solution quality (Liu et al., 2020). While a number of variations of LPA have been introduced, the original formulation remains the simplest and most efficient (Garza and Schaeffer, 2019).
A few open-source implementations and software packages have been developed for community detection utilizing the LPA. Fast Label Propagation Algorithm (FLPA) (Traag and Šubelj, 2023) is a variant of the LPA that employs a queue-based method, processing only vertices with recently updated neighborhoods. NetworKit (Staudt et al., 2016) is a software package built to analyze the structural attributes of graph datasets containing billions of connections. It integrates C++ kernels with a Python frontend, and features a parallel implementation of LPA. igraph (Csardi et al., 2006) is a similar package that is written in C, and offers Python, R, and Mathematica frontends. It has good adoption, and includes an LPA implementation.
Raghavan et al. (Raghavan et al., 2007) note that their Label Propagation Algorithm (LPA) for community detection can uncover internally-disconnected communities (Gregory, 2010). Genetic algorithms (Hesamipour et al., 2022) and expectation minimization/maximization algorithms (Ball et al., 2011; Hafez et al., 2014) can also detect such disconnected communities. To address this, they propose applying Breadth-First Search (BFS) to subnetworks in each individual group to separate the disconnected communities. This can be achieved with a time complexity of $O(M+N)$. Gregory (Gregory, 2010) extends LPA with the Community Overlap PRopagation Algorithm (COPRA), which removes nested communities and splits disconnected communities using a method similar to Raghavan et al. (Raghavan et al., 2007). Hafez et al. (Hafez et al., 2014) introduce a community detection algorithm using Bayesian Network and Expectation Minimization (BNEM), which includes a step to detect disconnected components within communities. If detected, disconnected components are assigned new community labels. This scenario arises when the network has more communities than the specified number $k$ in the algorithm. Hesamipour et al. (Hesamipour et al., 2022) propose a genetic algorithm for community detection, integrating similarity-based and modularity-based approaches. They use an MST-based representation to address issues like disconnected communities and ineffective mutations.
3. Preliminaries
Consider an undirected graph $G(V,E,w)$, where $V$ denotes the set of vertices, $E$ denotes the set of edges, and $w_{ij}=w_{ji}$ denotes the weight associated with each edge ($i$ and $j$ represent the vertices at the endpoints of each edge). In the case of unweighted graphs, we assume a unit weight for each edge, i.e., $w_{ij}=1$. The neighbors of a vertex $i$ are represented as $J_{i}=\{j\ |\ (i,j)\in E\}$, the weighted degree of each vertex is represented as $K_{i}=\sum_{j\in J_{i}}w_{ij}$, the total number of vertices is $N=|V|$, the total number of edges is $M=|E|$, and the sum of edge weights in the undirected graph is $m=\sum_{i,j\in V}w_{ij}/2$.
3.1. Community detection
Disjoint community detection involves assigning a community ID $c$ from set $\Gamma$ to each vertex $i\in V$, denoted by the mapping $C:V\rightarrow\Gamma$. Vertices belonging to community $c\in\Gamma$ are denoted as $V_{c}$, and the community of a vertex $i$ is represented by $C_{i}$. Further, we represent the neighbors of vertex $i$ within community $c$ as $J_{i\rightarrow c}=\{j\ |\ j\in J_{i}\ \text{and}\ C_{j}=c\}$, their corresponding edge weights summed as $K_{i\rightarrow c}=\sum_{j\in J_{i\rightarrow c}}w_{ij}$, the edge weight within community $c$ as $\sigma_{c}=\sum_{(i,j)\in E\ \text{and}\ C_{i}=C_{j}=c}w_{ij}$, and the total edge weight of community $c$ as $\Sigma_{c}=\sum_{(i,j)\in E\ \text{and}\ C_{i}=c}w_{ij}$ (Leskovec, 2021).
3.2. Modularity
Modularity is a metric that is used for assessing the quality of communities detected by heuristic-based algorithms. It quantifies the difference between the observed fraction of edges within communities and the expected fraction under random edge distribution. Modularity values range from $-0.5$ to $1$, where higher values indicate better community structure (Brandes et al., 2007). The modularity $Q$ is calculated using Equation 1, where $\delta$ represents the Kronecker delta function ($\delta(x,y)=1$ if $x=y$, 00 otherwise)))).
(1) $Q=\frac{1}{2m}\sum_{(i,j)\in E}\left[w_{ij}-\frac{K_{i}K_{j}}{2m}\right]\delta(C_{i},C_{j})=\sum_{c\in\Gamma}\left[\frac{\sigma_{c}}{2m}-\left(\frac{\Sigma_{c}}{2m}\right)^{2}\right]$
3.3. Label Progagation Algorithm (LPA)
LPA (Raghavan et al., 2007) is a widely used diffusion-based technique for identifying medium-quality communities in large networks. It is known for its simplicity, speed, and scalability. In LPA, each vertex $i$ initially possesses a unique label (community ID) $C_{i}$. During each iteration, vertices adopt the label with the highest interconnecting weight, as illustrated in Equation 2. This iterative process leads to a consensus among densely connected groups of vertices. The algorithm converges when at least $1-\tau$ fraction of vertices (where $\tau$ is the iteration tolerance parameter) maintain their community membership. LPA exhibits a time complexity of $O(L|E|)$ and a space complexity of $O(|V|+|E|)$, with $L$ representing the number of iterations performed (Raghavan et al., 2007). LPA has also been utilized in influence maximization, where the goal is to identify a small set of influential nodes in a network that can maximize the spread of information (Kumar et al., 2021).
(2) $C_{i}=\ \underset{c\ \in\ \Gamma}{\arg\max}{\sum_{j\in J_{i}\ |\ C_{j}=c}w_{ij}}$
(a) The initial community structure with $4$ communities
(b) After a few iterations, $C2$, $C3$, and $C4$ combine
(c) Following this, vertex $4$ transitions to $C3$
Figure 1. An example illustrating the potential of internally-disconnected communities with LPA (Sahu, 2024). Here, $C1$, $C2$, $C3$, and $C4$ represent four communities derived after one iteration of LPA, with vertices $1$ to $7$ belonging to community $C1$. Thicker lines are used here to indicate higher edge weights.
3.4. Possibility of Internally-disconnected communities with LPA
LPA has been noted to potentially identify internally disconnected communities (Raghavan et al., 2007). This is illustrated with an example in Figure 1. Initially, Figure 1(a) show the community structure after one iteration of LPA, with four communities labeled $C1$, $C2$, $C3$, $C4$, and vertices $1$ to $7$ grouped in $C1$ (i.e., vertices $1$ to $7$ have identified $C1$ as the most weighted label among their neighbors). Subsequent iterations, illustrated in Figure 1(b), result in relabeling of vertices in $C2$ and $C4$ into $C3$, due to strong internal connections, resulting in the combining of communities $C2$, $C3$, and $C4$ into $C3$. This results in $C3$ becoming the most weighted label for vertex $4$. Thus, $4$ changes its label to $C3$, resulting in internal disconnection of $C1$ as vertices $1$, $2$, $3$, $5$, $6$, and $7$ retain their locally popular labels. This phenomenon is common across various LPA implementations, including FLPA (Traag and Šubelj, 2023), igraph LPA (Csardi et al., 2006), and NetworKit LPA (Staudt et al., 2016), which identify up to $3.1\%$, $0.3\%$, and $14.5\%$ of communities that are internally-disconnected, respectively.
4. Approach
In the preceding section, we discussed the occurrence of internally-disconnected communities with LPA. However, this phenomenon is not exclusive to LPA and has been observed in various other community detection algorithms (Blondel et al., 2008; Hesamipour et al., 2022; Ball et al., 2011; Hafez et al., 2014). To address this issue, a commonly used approach involves splitting disconnected communities as a post-processing step (i.e., after the completion of all iterations of the community detection algorithm, when the community memberships of vertices have converged), known as Split Last (SL) using Breadth First Search (BFS) (Raghavan et al., 2007; Gregory, 2010; Hafez et al., 2014; Luecken, 2016; Wolf et al., 2019).
To split the internally-disconnected communities using the SL method, we investigate three different techniques: minimum-label-based Label Propagation (LP), minimum-label-based Label Propagation with Pruning (LPP), and Breadth First Search (BFS). We choose to explore LP and LPP techniques for splitting disconnected communities due to their inherent parallelizability.
Using the LP technique, each vertex in the graph is initially assigned a unique label corresponding to its vertex ID. Subsequently, in every iteration, each vertex within its assigned community selects the minimum label among its neighbors. This iterative process continues until convergence is achieved for all vertex labels. Consequently, each vertex acquires a unique label within its connected component and its respective community, thereby partitioning communities consisting of multiple connected components. On the other hand, the LPP technique integrates a pruning optimization step where only unprocessed vertices are processed. Once a vertex is processed, it is flagged as such and is only reactivated (or marked as unprocessed) if one of its neighbors alter their labels. The pseudocode for both the LP and LPP techniques is provided in Algorithm 1, with its explanation given in Section 4.1.
In contrast, the BFS technique for splitting internally disconnected communities involves selecting a random vertex within each community and identifying all reachable vertices from it as part of one subcommunity. If any vertices remain unvisited in the original community, another random vertex is selected from the remaining set, and the process repeats until all vertices within each community are visited. Consequently, the BFS technique facilitates the partitioning of connected components within each community. The pseudocode for this technique is presented in Algorithm 2, with detailed explanations provided in Section 4.2.
4.1. Explanation of LP/LPP technique
We now present the pseudocode for the parallel minimum-label-based Label Propagation (LP) and Label Propagation with Pruning (LPP) techniques, outlined in Algorithm 1, which are utilized for partitioning internally-disconnected communities. These techniques serve as a post-processing step (SL) at the end of the community detection algorithm. The function splitDisconnectedLp() accepts as input the graph $G(V,E)$ and the community memberships $C$ of vertices, and outputs the updated community memberships $C^{\prime}$ where all disconnected communities have been separated.
The algorithm commences in lines 8-11 by initializing the minimum labels $C^{\prime}$ of each vertex to their respective vertex IDs and marking all vertices as unprocessed. Lines 12-29 denote the iteration loop of the algorithm. Initially, the number of changed labels $\Delta N$ is set (line 13). Subsequently, an iteration of label propagation is performed (lines 14-27), followed by a convergence check (line 29). During each iteration, unprocessed vertices are handled concurrently. For each unprocessed vertex $i$, if the Label Propagation with Pruning (LPP) technique is used, it is marked as processed. The algorithm then identifies the minimum label $c^{\prime}_{min}$ within the community of vertex $i$ (lines 18-21) by considering only the labels of its neighbors belonging to the same community, and its own label. If $c^{\prime}_{min}$ differs from the current minimum label $C^{\prime}[i]$ (line 22), $C^{\prime}[i]$ is updated, $\Delta N$ is incremented to mark the change in label, and neighboring vertices of $i$ in the same community are marked as unprocessed (to be updated in the subsequent iteration). The label propagation loop (lines 12-29) continues until no further changes occur (line 29). Finally, the updated labels $C^{\prime}$, reflecting the community membership of each vertex with no internally-disconnected communities, are returned (line 30).
Algorithm 1 Split disconnected communities using (min) LP (Sahu, 2024).
1:$G(V,E)$: Input graph
2:$C$: Initial community membership/label of each vertex
3:$C^{\prime}$: Updated community membership/label of each vertex
4:$c^{\prime}_{min}$: Minimum connected label within the community
5:$\Delta N$: Number of changes in labels
6:
7:function splitDisconnectedLp($G,C$)
8:     $C^{\prime}\leftarrow\{\}$
9:     for all $i\in V$ in parallel do
10:         Mark $i$ as unprocessed
11:         $C^{\prime}[i]=i$
12:     loop
13:         $\Delta N\leftarrow 0$
14:         for all unprocessed $i\in V$ in parallel do
15:              if is SL-LPP or SP-LPP then
16:                  Mark $i$ as processed
[…]
18:              $c^{\prime}_{min}\leftarrow C^{\prime}[i]$
19:              for all $j\in G.out(i)$ do
20:                  if $C[j]=C[i]$ then
21:                       $c^{\prime}_{min}\leftarrow min(C^{\prime}[j],c^{\prime}_{min})$
22:              if $c^{\prime}_{min}=C^{\prime}[i]$ then continue
[…]
24:              $C^{\prime}[i]\leftarrow c^{\prime}_{min}$ ; $\Delta N\leftarrow\Delta N+1$
25:              if is SL-LPP or SP-LPP then
26:                  for all $j\in G.out(i)$ do
27:                       Mark $j$ as unprocessed if $C[j]=C[i]$
[…]
29:         if $\Delta N=0$ then break
30:     return $C^{\prime}$
4.2. Explanation of BFS technique
Following that, we move on to describe the pseudocode of the parallel Breadth First Search (BFS) technique, as outlined in Algorithm 2, designed for the partitioning of disconnected communities. Similar to the LP/LPP techniques, this approach can be utilized as a post-processing step (SL) at the conclusion of the process. The function splitDisconnectedBfs() receives the input graph $G(V,E)$ and the community membership $C$ of each vertex, and yields the updated community membership $C^{\prime}$ of each vertex where all the disconnected communities have been separated.
Initially, between lines 10 and 12, the flag vector $vis$, indicating visited vertices, is initialized, and the labels $C^{\prime}$ for each vertex are set to their respective vertex IDs. Subsequently, multiple thread concurrently process each vertex $i$ in the graph $G$ (lines 14 to 19). If the community $c$ of vertex $i$ is not present in the work-list $work_{t}$ of the current thread $t$, or if vertex $i$ has already been visited, the thread proceeds to the next iteration (line 16). Conversely, if community $c$ is in the work-list $work_{t}$ of the current thread $t$ and vertex $i$ has not been visited, a Breadth-First Search (BFS) is executed from vertex $i$ to explore vertices within the same community. This BFS employs lambda functions $f_{if}$ to selectively perform BFS on vertex $j$ if it belongs to the same community, and $f_{do}$ to update the label of visited vertices after each vertex is explored during BFS (line 19). After processing all vertices, the threads synchronize, and the updated labels $C^{\prime}$, denoting the community membership of each vertex with no disconnected communities, are returned (line 20). It is worth noting that the work-list $work_{t}$ for each thread identified by $t$ is defined as a set containing communities $[t\chi,\ t(\chi+1))\ \cup\ [T\chi+t\chi,\ T\chi+t(\chi+1))\ \cup\ \ldots$, where $\chi$ represents the chunk size and $T$ denotes the total number of threads used. In our implementation, a chunk size of $\chi=1024$ is utilized for efficient parallel processing. This choice optimizes workload distribution among threads and enhances overall performance during community detection tasks.
Algorithm 2 Split disconnected communities using BFS (Sahu, 2024).
1:$G(V,E)$: Input graph
2:$C$: Initial community membership/label of each vertex
3:$C^{\prime}$: Updated community membership/label of each vertex
4:$f_{if}$: Perform BFS to vertex $j$ if condition satisfied
5:$f_{do}$: Perform operation after each vertex is visited
6:$vis$: Visited flag for each vertex
7:$work_{t}$: Work-list of current thread
8:
9:function splitDisconnectedBfs($G,C$)
10:     $C^{\prime}\leftarrow\{\}$ ; $vis\leftarrow\{\}$
11:     for all $i\in V$ in parallel do
12:         $C^{\prime}[i]=i$
13:     for all threads do
14:         for all $i\in V$ do
15:              $c^{\prime}\leftarrow C^{\prime}[i]$
16:              if $c\notin work_{t}$ or $vis[i]$ then continue
17:              $f_{if}\leftarrow(j)\implies C[j]=C[j]$
18:              $f_{do}\leftarrow(j)\implies C^{\prime}[j]\leftarrow c^{\prime}$
19:              $bfsVisitForEach(vis,G,i,f_{if},f_{do})$
20:     return $C^{\prime}$
Figure 2 depicts an example of the Breadth-First Search (BFS) technique. Initially, as shown in Figure 2(a), two communities, $C1$ and $C2$, are formed after running a few iterations of LPA. Here, $C1$ experiences internal disconnection due to the vertex $4$ selecting $C2$ as its best community, mirroring the scenario illustrated in Figure 1(c). Subsequently, employing the BFS technique, a thread selects a random vertex within community $C1$, such as $2$, and assigns the label $2$ to all vertices reachable within $C1$ from vertex $2$, marking them as visited (Figure 2(b)). Following this, as illustrated in Figure 2(c), the same thread randomly selects an unvisited vertex within community $C1$, for instance, $7$, and assigns the label $7$ to all vertices reachable within $C1$ from vertex $7$, marking them as visited. A similar procedure is carried out within community $C2$. As a result, all vertices are visited, and the labels assigned to them signify the updated community membership of each vertex without disconnected communities. It’s essential to note that each thread possesses a distinct work-list, ensuring that concurrent BFS operations within the same community are avoided.
We also attempted another parallel BFS-based algorithms to split disconnected communities. It begins by initializing the visited flags and BFS frontiers, using OpenMP to parallelize across vertices. Then, for each community, it selects a starting vertex and marks it as visited. Using a BFS approach, it iteratively explores neighbors within each community, updating each unvisited vertex’s community label, and marking it as visited if it shares the same community. The algorithm continues this process in a loop, advancing the BFS frontier until no new vertices are marked in a given iteration, at which point a reinitialization is triggered for untouched vertices in each community. This procedure is repeated until all vertices are processed. However, this algorithm offered lower performance than our BFS-based algorithm discussed above.
(a) Community $C1$ is internally disconnected
(b) Post BFS relabeling from vertex $2$
(c) Post BFS relabeling from vertex $7$
Figure 2. An example demonstrating the BFS technique for splitting disconnected communities (Sahu, 2024). Initially, two communities, $C1$ and $C2$, are depicted, where $C1$ is internally disconnected because vertex $4$ has joined $C2$. The BFS technique randomly selects vertices within each community and assigns the same label to reachable vertices (indicated with a new community ID).
4.3. Our GSL-LPA algorithm
To evaluate the effectiveness of the Split Last (SL) approach in conjunction with the minimum-label-based Label Propagation (LP), minimum-label-based Label Propagation with Pruning (LPP), and Breadth-First Search (BFS) techniques for splitting disconnected communities within the framework of LPA, we employ GVE-LPA (Sahu, 2023b), our parallel implementation of LPA.
(a) Relative runtime using Split Last (SL) approach, based on three different techniques, for splitting disconnected communities with Parallel LPA
(b) Modularity of communities obtained using Split Last (SL) approach for splitting disconnected communities with Parallel LPA
(c) Fraction of disconnected communities (logarithmic scale) using Split Last (SL) approach for splitting disconnected communities with Parallel LPA
Figure 3. Mean relative runtime, modularity, and fraction of disconnected communities (log-scale) using the Split Last (SL) approach for addressing disconnected communities with Parallel LPA (Sahu, 2023c) across all graphs in the dataset. The SL approach utilizes minimum label-based Label Propagation (LP), Label Propagation with Pruning (LPP), or Breadth First Search (BFS) techniques for splitting disconnected communities.
4.3.1. Determining suitable technique for splitting disconnected communities
To evaluate the optimal technique for partitioning disconnected communities, we investigate the SL approach with LP, LPP, and BFS techniques. Figures 3(a), 3(b), and 3(c) display the mean relative runtime, modularity, and fractions of disconnected communities for SL-LP, SL-LPP, SL-BFS, and the default approaches. The SL approach yields non-disconnected communities, as depicted in Figure 3(c). Additionally, Figure 3(b) shows that the modularity of communities obtained through the SL approach exceeds that of the default approach. Figure 3(a) illustrates that SL-BFS, i.e., the SL approach with BFS technique, exhibits superior performance. Thus, employing BFS for splitting disconnected communities as a post-processing step (SL) of LPA is preferable.
Algorithm 3 GSL-LPA: Our Parallel LPA which identifies communities that are not internally disconnected.
1:$G(V,E)$: Input graph
2:$C$: Community membership of each vertex
3:$H_{t}$: Collision-free per-thread hashtable
4:$l_{i}$: Number of iterations performed
5:$\tau$: Per iteration tolerance
6:
7:function lpa($G$)
8:     Vertex membership: $C\leftarrow[0..|V|)$
9:     Mark all vertices in $G$ as unprocessed
[…]
11:         $\Delta N\leftarrow lpaMove(G,C)$
12:         if $\Delta N/N\leq\tau$ then break $\triangleright$ Converged?
13:     $C\leftarrow splitDisconnectedBfs(G,C)$triangleright Algorithm 2
14:     return $C$
15:
16:function lpaMove($G,C$)
17:     Changed vertices: $\Delta N\leftarrow 0$
18:     for all unprocessed $i\in V$ in parallel do
19:         Mark $i$ as processed (prune)
20:         $H_{t}\leftarrow scanCommunities(\{\},G,C,i)$
[…]
22:         $c^{*}\leftarrow$ Most weighted label to $i$ in $G$
23:         if $c^{*}=C[i]$ then continue
24:         $C[i]\leftarrow c^{*}$ ; $\Delta N\leftarrow\Delta N+1$
25:         Mark neighbors of $i$ as unprocessed
26:     return $\Delta N$
27:
28:function scanCommunities($H_{t},G,C,i$)
29:     for all $(j,w)\in G.edges(i)$ do
30:         if $i\neq j$ then $H_{t}[C[j]]\leftarrow H_{t}[C[j]]+w$
31:     return $H_{t}$
4.3.2. Explanation of the algorithm
We denote GVE-LPA utilizing the Split Last (SL) approach with Breadth-First Search (SL-BFS) technique to deal with the internally-disconnected communities as GSL-LPA. The core procedure of GSL-LPA, encapsulated in the lpa() function, is presented in lines 7-14 of Algorithm 3. This function takes a graph $G(V,E)$ as input and produces the community membership (or label) $C$ for each vertex in the graph, ensuring that none of the returned communities are internally disconnected.
In lines 8-9, we start by initializing the labels $C$ for each vertex in $G$, and marking all vertices in the graph as unprocessed. We then iterate to propagate labels based on the weighted influence of neighboring vertices. We limit the number of iterations to $MAX\_ITERATIONS$ (lines 10-12). Within each iteration, we utilize the lpaMove() function for label propagation and track the count of nodes with updated labels $\Delta N$ (line 11). If the ratio of $\Delta N$ to the total number of nodes $N$ falls within the specified iteration tolerance $\tau$, convergence is achieved, and the loop terminates (line 12). After completing all iterations, the algorithm enters the splitting phase to separate internally-disconnected communities in $C$. This is accomplished using the parallel BFS technique in line 13, employing the splitCommunitiesBfs() function (Algorithm 2). Finally, we return the updated labels $C$ (line 14).
In lines 16-26, the function lpaMove() operates by concurrently iterating over unprocessed vertices. For each unprocessed vertex $i$ in the graph $G$, it marks $i$ as processed (vertex pruning optimization, line 19), computes the total edge weight of connected labels in a per-thread hashtable $H_{t}$ using the scanCommunities() function (line 20), and selects the most weighted label $c^{*}$ (line 22). If $c^{*}$ differs from the current label of $i$, the label of $i$ is updated, the count of changed vertices $\Delta N$ is incremented, and the neighbors of $i$ are marked as unprocessed for the next iteration (lines 24-25). Upon having processed all the vertices, the function returns the total number of vertices with updated labels $\Delta N$ (line 26). The scanCommunities() function, given in lines 28-31, iterates over the neighbors of the current vertex $i$, excluding itself, and computes the total edge weight linked to each label in the hashtable $H_{t}$.
5. Evaluation
5.1. Experimental Setup
5.1.1. System used
We employ a server consisting of two Intel Xeon Gold 6226R processors, each featuring $16$ cores running at $2.90$ GHz. Each core includes a $1$ MB L1 cache, a $16$ MB L2 cache, and shares a $22$ MB L3 cache. The system is equipped with $376$ GB of RAM and runs CentOS Stream 8.
5.1.2. Configuration
We utilize 32-bit integers for representing vertex IDs and 32-bit floats for edge weights, while computations and hashtable values are based on 64-bit floats. We use $64$ threads by default, which to correspond to the available cores on the system. Compilation is conducted using GCC 8.5 and OpenMP 4.5.
5.1.3. Dataset
The graphs employed in our experiments are detailed in Table 1, obtained from the SuiteSparse Matrix Collection (Kolodziej et al., 2019). These graphs encompass a range of $3.07$ to $214$ million vertices and $25.4$ million to $3.80$ billion edges. We make sure that the edges are both undirected and weighted, with a default weight set to $1$.
Table 1. List of $13$ graphs sourced from the SuiteSparse Matrix Collection (Kolodziej et al., 2019) (directed graphs are marked with *). Here, $|V|$ is the number of vertices, $|E|$ is the number of edges (after adding reverse edges), $D_{avg}$ is the average degree, and $|\Gamma|$ is the number of communities identified with GSL-LPA.
Graph $|V|$ $|E|$ $D_{avg}$ $|\Gamma|$
Web Graphs (LAW)
indochina-2004 7.41M 341M 41.0 160K
uk-2002 18.5M 567M 16.1 414K
arabic-2005 22.7M 1.21B 28.2 268K
uk-2005 39.5M 1.73B 23.7 744K
webbase-2001 118M 1.89B 8.6 6.70M
it-2004 41.3M 2.19B 27.9 694K
sk-2005 50.6M 3.80B 38.5 433K
Social Networks (SNAP)
com-LiveJournal 4.00M 69.4M 17.4 6.91K
com-Orkut 3.07M 234M 76.2 192
Road Networks (DIMACS10)
asia_osm 12.0M 25.4M 2.1 278K
europe_osm 50.9M 108M 2.1 1.52M
Protein k-mer Graphs (GenBank)
kmer_A2a 171M 361M 2.1 40.1M
kmer_V1r 214M 465M 2.2 48.7M
(a) Runtime in seconds (logarithmic scale) with FLPA, igraph LPA, NetworKit LPA, and GSL-LPA
(b) Speedup (logarithmic scale) of GSL-LPA with respect to FLPA, igraph LPA, NetworKit LPA.
(c) Modularity of communities obtained with FLPA, igraph LPA, NetworKit LPA, and GSL-LPA.
(d) Fraction of disconnected communities (logarithmic scale) with FLPA, igraph LPA, NetworKit LPA, and GSL-LPA.
Figure 4. Runtime in seconds (log-scale), speedup (log-scale), modularity, and fraction of disconnected communities (log-scale) compared across FLPA, igraph LPA, NetworKit LPA, and GSL-LPA for each graph in the dataset. igraph LPA fails to execute on kmer_A2a and kmer_V1r graphs, and thus its results are excluded.
5.2. Comparing Performance of GSL-LPA
We now compare the performance of GSL-LPA with igraph LPA (Csardi et al., 2006), FLPA (Traag and Šubelj, 2023), and NetworKit LPA (Staudt et al., 2016). Both igraph LPA and FLPA are implemented sequentially, while NetworKit LPA is parallelized. For FLPA, we checkout the branch featuring the modified version of igraph_community_label_propagation() function, and update the label propagation example in C to use the IGRAPH_LPA_FAST variant, and enable loading the input graph from a file. The runtime of igraph_community_label_propagation() is measured using gettimeofday(). For igraph LPA, we adopt a similar methodology as FLPA, but utilize the code from the master branch of igraph. For NetworKit, we use a Python script to run PLP (Parallel Label Propagation) algorithm, and record the runtime with getTiming(). The runtime of NetworKit LPA is measured five times for each graph to obtain an average. Due to extensive duration of igraph LPA runs, we perform only a single run per graph. Additionally, we record the modularity of communities identified by each implementation.
Figure 4(a) depicts the runtimes of FLPA, igraph LPA, NetworKit LPA, and GVE-LPA for each graph in the dataset. Meanwhile, Figure 4(b) illustrates the speedup of GSL-LPA compared to each of the aforementioned implementations of LPA. igraph LPA encounters runtime issues on protein k-mer graphs (kmer_A2a and kmer_V1r), hence its results for these graphs are omitted.
GSL-LPA demonstrates average speedups of $\times5$, 10,300×10,300$\times$10 , 300 ×, and 5.8×5.8$\times$5.8 × over FLPA, igraph LPA, and NetworKit LPA respectively. Specifically, on the sk-2005 graph, GSL-LPA identifies communities within $4.5$ seconds, achieving a processing rate of $844$ million edges/s. Furthermore, Figure 4(c) presents the modularity of communities obtained by each implementation. On average, GSL-LPA achieves modularity that is $7.1\%$ and $0.7\%$ higher than FLPA (particularly on road networks and protein k-mer graphs) and igraph LPA respectively. Additionally, it records a $3.6\%$ lower modularity compared to NetworKit LPA (especially on protein k-mer graphs). Finally, Figure 4(d) illustrates the fraction of disconnected communities obtained by each implementation. The absence of bars indicates the absence of disconnected communities. Communities identified by the GSL-LPA exhibit no disconnected communities. However, on average, FLPA, igraph LPA, and NetworKit LPA exhibit fractions of disconnected communities amounting to $2.3\times 10^{-2}$, $1.2\times 10^{-3}$, and $1.8\times 10^{-2}$, respectively, particularly noticeable on web graphs and social networks. Figure 7 depicts the comparison of GVE-LPA and GSL-LPA. This comparison is explained in detail in Section A.2.
Figure 5. Phase split of GSL-LPA for each graph in the dataset.
Figure 6. Overall speedup of GSL-LPA, including its label propagation and splitting phases, with increasing number of threads (in powers of 2).
5.3. Performance Analysis of GSL-LPA
Next, we proceed with an analysis of GSL-LPA’s performance. The phase-wise breakdown of GSL-Leiden is depicted in Figure 5. This illustration shows that GSL-LPA allocates a significant portion of its runtime to the splitting phase, particularly noticeable on social networks, whereas it predominantly emphasizes the label propagation phase on web graphs. On average, GSL-LPA dedicates $47\%$ of its runtime to the label propagation phase and $53\%$ to the splitting phase. We plan to address the optimization of the splitting phase, i.e., Algorithm 2, in future work.
5.4. Strong Scaling of GSL-LPA
Finally, we evaluate the strong scaling performance of GSL-LPA. This involves varying the number of threads from $1$ to $64$ in powers of $2$, for each input graph, and measuring the time taken for GSL-LPA to identify communities. Figure 6 illustrates the average strong scaling of GSL-LPA, including the scaling of its individual phases: the label propagation phase and the splitting phase. At 32 threads, GSL-LPA achieves a speedup of 10.7×10.7$\times$10.7 × compared to single-threaded execution, indicating a performance increase of 1.6×1.6$\times$1.6 × for each doubling of threads. The scalability is constrained by the splitting phase, which only provides a speedup of 6.2×6.2$\times$6.2 × with 32 threads, while the label propagation phase offers a speedup of 14.5×14.5$\times$14.5 × with the same number of threads. At 64 threads, GSL-LPA is affected by NUMA effects and offers a speedup of 11.5×11.5$\times$11.5 ×.
6. Conclusion
In this report, we explored the issue of internally-disconnected communities arising from the Label Propagation Algorithm (LPA). Our experimental findings showed that FLPA (Traag and Šubelj, 2023), igraph LPA (Csardi et al., 2006), and NetworKit LPA (Staudt et al., 2016) identified communities with this issue in up to $3.1\%$, $0.3\%$, and $14.5\%$ of cases respectively. To address this, we introduced GSL-LPA, derived from our parallelization of LPA (GVE-LPA) (Sahu, 2023c), which mitigates this problem by splitting disconnected communities as a post-processing step, akin to previous approaches. Our experiments conducted on a system equipped with two 16-core Intel Xeon Gold 6226R processors demonstrated that GSL-LPA not only resolves this issue but also outperforms FLPA, igraph LPA, and NetworKit LPA by $\times5$, 10,300×10,300$\times$10 , 300 ×, and 5.8×5.8$\times$5.8 × respectively. Moreover, GSL-LPA achieves a processing rate of $844M$ edges/s on a $3.8B$ edge graph. Additionally, GSL-LPA exhibits performance improvement at a rate of 1.6×1.6$\times$1.6 × for every doubling of threads.
Acknowledgements.
I would like to thank Prof. Kishore Kothapalli and Prof. Dip Sankar Banerjee for their support.
Appendix A Appendix
A.1. Finding disconnected communities
We present a parallel algorithm aimed at identifying disconnected communities using the original graph and vertex community memberships as input. The fundamental principle involves evaluating each community’s size, selecting a representative vertex, traversing within the community while avoiding neighboring communities, and flagging a community as disconnected if all its vertices are unreachable. We explore four distinct approaches, characterized by their utilization of parallel Depth-First Search (DFS) or Breadth-First Search (BFS), and whether per-thread or shared visited flags are employed. In the case of shared visited flags, each thread scans all vertices but exclusively processes its assigned community based on the community ID. Our results demonstrate that employing parallel BFS traversal with a shared flag vector yields the most efficient outcomes. Due to the deterministic nature of this algorithm, all approaches produce identical results. Algorithm 4 presents the pseudocode for this approach, where the disconnectedCommunities() function takes the input graph $G$ and community membership $C$, returning the disconnected flag $D$ for each community.
Algorithm 4 Finding disconnected communities in parallel (Sahu, 2023a).
1:$G(V,E)$: Input graph
2:$C$: Community membership of each vertex
3:$D$: Disconnected flag for each community
4:$S$: Size of each community
5:$f_{if}$: Perform BFS to vertex $j$ if condition satisfied
6:$f_{do}$: Perform operation after each vertex is visited
7:$reached$: Number of vertices reachable from $i$ to $i$’s community
8:$vis$: Visited flag for each vertex
9:$work_{t}$: Work-list of current thread
10:
11:function disconnectedCommunities($G,C$)
12:     $D\leftarrow\{\}$ ; $vis\leftarrow\{\}$
13:     $S\leftarrow communitySizes(G,C)$
14:     for all threads in parallel do
15:         for all $i\in V$ do
16:              $c\leftarrow C[i]$ ; $reached\leftarrow 0$
[…]
19:              if $S[c]=0$ or $c\notin work_{t}$ then continue
20:              $f_{if}\leftarrow(j)\implies C[j]=c$
21:              $f_{do}\leftarrow(j)\implies reached\leftarrow reached+1$
22:              $bfsVisitForEach(vis,G,i,f_{if},f_{do})$
23:              if $reached<S[c]$ then $D[c]\leftarrow 1$
24:              $S[c]\leftarrow 0$
25:     return $D$
Let us now explain Algorithm 4. Initially, in line 12, we set up the disconnected community flag $D$ and the visited vertices flags $vis$. Line 13 calculates the size of each community $S$ concurrently using the communitySizes() function. Subsequently, each thread processes every vertex $i$ in the graph $G$ concurrently (lines 15-24). In line 16, we ascertain the community membership of $i$ ($c$) and initialize the count of reached vertices from $i$ to 00. If community $c$ is either empty or not in the work-list of the current thread $work_{t}$, the thread proceeds to the next iteration (line 19). However, if community $c$ is non-empty and in the work-list of the current thread $work_{t}$, we execute BFS from vertex $i$ to explore vertices within the same community. This involves utilizing lambda functions $f_{if}$ to conditionally execute BFS to vertex $j$ if it belongs to the same community and $f_{do}$ to update the count of reached vertices after each vertex is visited during BFS (line 22). If the number of vertices $reached$ during BFS is less than the community size $S[c]$, we flag community $c$ as disconnected (line 23). Finally, we update the size of the community $S[c]$ to 00, indicating that the community has been processed (line 24). It’s worth noting that the work-list $work_{t}$ for each thread with ID $t$ is defined as a set containing communities $[t\chi,\ t(\chi+1))\ \cup\ [T\chi+t\chi,\ T\chi+t(\chi+1))\ \cup\ \ldots$, where $\chi$ is the chunk size and $T$ is the number of threads. In our implementation, we use a chunk size of $\chi=1024$.
A.2. Additional Performance comparison
Next, we proceed with the comparison of performance between GSL-LPA and GVE-LPA (Sahu, 2023c). Following a similar methodology as earlier, we run both implementations five times for every graph in the dataset to reduce measurement noise, and present the averages in Figures 7(a), 7(b), 7(c), and 7(d).
Figure 7(a) illustrates the runtimes of GVE-LPA and GSL-LPA on each graph in the dataset. On average, GSL-LPA demonstrates a runtime approximately $125\%$ longer than GVE-LPA. This additional computational time represents a trade-off made to ensure the absence of internally disconnected communities. To minimize the runtime for GSL-LPA, it is necessary to optimize Algorithm 2. We plan to address this in the future.
Figure 7(c) showcases the modularity of communities obtained by each implementation. On average, GSL-LPA achieves a $0.4\%$ higher modularity compared to GVE-LPA. Finally, Figure 7(d) presents the fraction of internally disconnected communities identified by each implementation. Communities obtained with GSL-LPA exhibit no disconnected communities, while those identified with GVE-LPA feature an average of $6.6\%$ disconnected communities.
(a) Runtime in seconds (logarithmic scale) with GVE-LPA and GSL-LPA
(b) Speedup of GVE-LPA with respect to GSL-LPA.
(c) Modularity of communities obtained with GVE-LPA and GSL-LPA.
(d) Fraction of disconnected communities (logarithmic scale) with GVE-LPA and GSL-LPA.
Figure 7. Runtime in seconds (log-scale), speedup, modularity, and fraction of disconnected communities (log-scale) with GVE-LPA (Sahu, 2023c) and GSL-LPA for each graph in the dataset.